A 0.025g sample of aspirin is saponified, treated with excess Fe3+(aq) and diluted to a final volume of 1.00liter. In the spectrometer an absorbance that corresponds to a concentration of 1.35x10-4M is recorded. What % of the sample's mass is due to aspirin?


1
Expert's answer
2021-02-22T05:55:21-0500

molarity = no . Of moles / volume

1.35×10-4 = moles / 1 L

moles = 1.35×10-4

Mass of Fe3+= 1.35×10-4 × 55.84 = 0.0075 g

Total mass = Mass of aspirin + Mass of Fe3+

= 0.025 + 0.0075 = 0.0325

% Mass = "\frac{0.025}{0.0325}" ×100

% Mass = 76.9 % of the sample's mass is due to aspirin